OpenStudy (calculusxy):

MEDAL AND FAN FOR WHOEVER HELPS ME SOLVE SOME QUESTIONS!!!!!!!!! Ella spent her whole allowance of $2.00, plus the 16¢ she had left over from last week, on bubble gum. If the pieces of gum had been a penny cheaper, she would have received three more pieces than she did. How many pieces did she actually buy? (A) 8 (B) 9 (C) 16 (D) 24 (E) 27

OpenStudy (calculusxy):

@goformit100

OpenStudy (anonymous):

let cost of bubblegum =x cost of bubblegum if got 1 penny cheaper = x-1 let it buy n bub for 2$ 16 cent or 216 cent let it n bubble gum for x each -> nx=216 1 let it buy n+3 bubble for x-1 ->(n+3)(x-1)=216 2 now solve 1 and 2

OpenStudy (calculusxy):

The equation that you have given me, is something that I can't solve. I need help with solving that equation.

OpenStudy (anonymous):

nx +3x-n-3 =216 x=216/n =216+3*216/n-n-3=216 3*216=n(n+3) 648=n^2+3n n= 24,-27 neglecting -27 n=24 bubblegums

OpenStudy (calculusxy):

@harpreetsk Is there any easier way to explain on what you did? Because I simply dont get it.

OpenStudy (anonymous):

basicly i simlified equation 1 and 2 first as xn=216 so x=216/n .........p and simplified other by multiplying it to get nx +3x-n-3 =216 .....q then substituted pp in q and simlyfied it

OpenStudy (anonymous):

no convert cost into penny as the changes are given in penny